OpenStudy (anonymous):

A line has equation y= -3/7x+5. Determine the slope angle of the line, to the nearest degree.

OpenStudy (anonymous):

\(\tan(\theta)=-\frac{3}{7}\) so \(\theta=\tan^{-1}(-\frac{3}{7})\) then a calculator

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=arctan%28-3%2F7%29 yw